5) 13,472.52 Question 7 (1 point) The temperature in Grand Haven in August (fahrenheit) for 10 days was 768:791:86.2 69.7 73.4; 77.5; 774; 80.2; 83.3, 88.2. Calculate the coefficient of variation of the dataset. 1) 444.472 2) -73.567 3) 11.227 4) 84.793 5) 0.071 average speed of the cars was 70.5 and the standard devilation was 4.86. However, since th radar gun was biased. 4 is added to every observation in the dataset. What will the new mean Question 8 (1 point) Suppose that a police officer measured the speed of cars on a highway for an hour. The
Answer #1

7)

(x-x) 2.380 S.No (X-X) 5.6644 0.0064 49.2804 89.8704 33.4084 2.8224 3.1684 1 76.80 79.10 86.20 69.70 73.40 77.50 77.40 80.20 83.30 88.20 791.8 2 7.020 9.480 -5.780 -1.680 -1.780 1.020 4.120 9.020 4 16.9744 81.3604 283.5960 31.51 5.61 10 2x 79.1800 -2(x-X)/(n-1)-

from above coefficient of variation =(std deviation/mean)*100 =(5.61/79.18)=0.071

8)

new mean =70.5+4=74.5